Geometry. A square has sides of length 3x - 2 cm. Express the area of the square as a polynomial.
I have no idea. Please help.
Thanks.

Answer by sdmmadam@yahoo.com(530)

A square has sides of length 3x - 2 cm.
Express the area of the square as a polynomial.
A square is a quadrilateral with all its sides equal and each angle a right angle that is 90 degrees.
If a square has side y cms, then its area is given by y^2 sq cms
Therefore given that y = (3x-2)cms
The area of the given square is (3x-2)^2 sq cms
= (3x)^2 + 2(3x)X(-2) + (-2)^2 [using the formula (a+b)^2 =a^2 +2ab + b^2 ]
= 9x^2 -12x + 4
Answer: (9x^2 -12x + 4) sq cms
Note: You may write (3x-2)^2 as
(3x-2)(3x-2) = (3x)(3x-2)-2(3x-2)=(3x)^2-6x-6x +4 = 9x^2 -12x +4